40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Mar 2 19:52:46 2021, gautam, Update, LSC, REFL55 demod board rework 
    Reply  Wed Mar 3 22:13:05 2021, gautam, Update, LSC, REFL55 demod board rework 
       Reply  Thu Mar 4 23:16:08 2021, Koji, Update, LSC, REFL55 demod board rework D990511-00_REFL55.pdfP_20210304_215602.jpgP_20210304_222400.jpgP_20210304_222412.jpg20210304234400_IMG_0526.jpg
          Reply  Fri Mar 5 13:53:57 2021, gautam, Update, LSC, REFL55 demod board rework 
             Reply  Fri Mar 5 15:31:23 2021, Koji, Update, LSC, REFL55 demod board rework 
                Reply  Fri Mar 5 16:24:24 2021, gautam, Update, LSC, REFL55 demod board re-installed in 1Y2 REFL55.pdfdemodNoise.pdf
                   Reply  Fri Mar 5 22:25:13 2021, gautam, Update, LSC, PRMI 1f SB locking recovered PRMI1f_noArmssensMat.pdf
Message ID: 15859     Entry time: Wed Mar 3 22:13:05 2021     In reply to: 15855     Reply to this: 15864
Author: gautam 
Type: Update 
Category: LSC 
Subject: REFL55 demod board rework 

After this work, I measured that the orthogonality was poor. I confirmed on the bench that the PQW-2-90 was busted, pin 2 (0 degree output) showed a sensible signal half of the input, but pin 6 had far too small an output and the phase difference was more like 45 degrees and not 90 degrees. I can't find any spares of this part in the lab - however, we do have the equivalent part used in the aLIGO demodulator. Koji has kindly agreed to do the replacement (it requires a bit of jumper wiring action because the pin mapping between the two parts isn't exactly identical - in fact, the circuit schematic uses a transformer to do the splitting, but at some unknown point in time, the change to the minicircuits part was made. Anyway, until this is restored, I defer the PRMI sideband locking.

Quote:

There were multiple problems with the REFL55 demod board. I fixed them and re-installed the board. The TFs and noise measured on the bench now look more like what is expected from a noise model. The noise in-situ also looked good. After this work, my settings for the PRMI sideband lock don't work anymore so I probably have to tweak things a bit, will look into it tomorrow.

ELOG V3.1.3-